I'm having trouble with this problem and I need a little help with it:
Investing $1000 in two investments yielding you a one-year return of $75. If one of the investments was a bond fund with a guranteed 4% rate of return, and you invested an equal amount in both, what was the rate of return in the second investment?
Here are the answers I have to choose from:
11%
10.5%
3.5%
12%
and I'v been trying to solve it with this formula :1000(1+r)=?

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

.04*500+500x=75
20+500x=75
500x=55
x=.11 or 11%

Answer by stanbon(75887)

Investing $1000 in two investments yielding you a one-year return of $75. If one of the investments was a bond fund with a guranteed 4% rate of return, and you invested an equal amount in both, what was the rate of return in the second investment?
--------------------------
Equation:
interest + interest = interest



0.04(500) + r(500) = 75
20 + 500r = 75
500r = 55
r = 55/500
r = 0.11
The rate you want is 11%
==================================
Cheers,
Stan H.
=====================================
Here are the answers I have to choose from:
11%
10.5%
3.5%
12%